OpenStudy (rebeccapink6398):

WILL MEDAL! Which decimal is equivalent to 5 9/20 A. 5.95 B. 5.45 C. 5.92 D. 5.9

OpenStudy (rebeccapink6398):

@31356 PLEASE HELP :)

OpenStudy (anonymous):

I think it might be B. if not I am sorry!D:

OpenStudy (anonymous):

5.45

OpenStudy (anonymous):

Yeaaaa I got it right X"D

OpenStudy (anonymous):

what you do is divide 9 by 20, which gives you .45

OpenStudy (anonymous):

Then add 5 back in!

OpenStudy (anonymous):

An easy way to do this without a calculator is to find what multiplied by 20 is 100. The answer is 5. So multiply 5 times your numerator as well, and you get 45. 45/100. From there you can easily see that 45/100 = 0.45. The add the 5 to that. 5.45.

OpenStudy (rebeccapink6398):

Thank you :)

OpenStudy (anonymous):

Of course, sometimes it's not easy if you can't easily identify a multiple of your denominator that will make it equal 100. But for a denominator like 20, it's easy to see that 5*20 = 100.

OpenStudy (anonymous):

I should say scalar, not multiple...

OpenStudy (rebeccapink6398):

Could you also help me with this one? Becca's family has 4 gallons of milk. They drink 5/8 gallon per day. How long will their milk last? A. 4 5/8days B. 2 1/2 days C. 4 days D. 6 2/5days

OpenStudy (rebeccapink6398):

ps sorry for all the double spacing

OpenStudy (anonymous):



OpenStudy (anonymous):

So do the math, (4) / (5/8) = ?

OpenStudy (rebeccapink6398):

????I I don't get it????

OpenStudy (anonymous):

Alright, so this problem involves something called "dimensional analysis." It is where you use conversion factors to convert from one unit to another. In this case, you are converting 4 gallons of milk into number of days the milk will last. You start at the left side with 4 gallons. The conversion is that they drink (5/8) gallons in 1 day. You set up your dimensional analysis so that your units cancel. I.e., you if you start with gallons, you need to multiply it by the conversion factor that has gallons in the denominator, so when you multiply across, those units cancel, and you are left with days.

OpenStudy (rebeccapink6398):

WOW!!! now I really don't get it

OpenStudy (anonymous):

\[4 gallons \times \frac{1 day}{\frac{5}{8}gallons} = \frac{4gallons\times1day}{\frac{5}{8}gallons} = \frac{4\times1day}{\frac{5}{8}}\]

OpenStudy (anonymous):

Can you make sense of that sequence of steps?

OpenStudy (rebeccapink6398):

no :( im not tring to get the answer out of you or anything I just don't get it :(

OpenStudy (anonymous):

\[= \frac{4 days}{\frac{5}{8}} = \frac{4days}{1}\times\frac{8}{5} = \frac{32days}{5} = \frac{32}{5}days\]

OpenStudy (anonymous):

That's okay. What grade level is this course? I think I am explaining it as if it were a college chem class. Perhaps someone else can put it in terms you will understand better than I can. But your answer is 32/5.

OpenStudy (anonymous):

Try to make sense of the steps I wrote from left to right.

OpenStudy (rebeccapink6398):

thank you :) and its 6th

OpenStudy (anonymous):

Oh wow, well you're doing great! :)

OpenStudy (rebeccapink6398):

(ps its not one of the options :( )

OpenStudy (anonymous):

is 6.4 an option?

OpenStudy (rebeccapink6398):

6 2/5 ps 6 is big number and 2/5

OpenStudy (anonymous):

I see. So 32 / 5 = 6 and 2/5 You see that 5 goes into 32 6 times, right? and you are left with 2/5

OpenStudy (rebeccapink6398):